OpenStudy (anonymous):

PLEASSSEEEE help 1/2log(low8)x=3log(low8)3

OpenStudy (anonymous):

kk hold on a second,

OpenStudy (anonymous):

so ur trying to solve for x?

OpenStudy (anonymous):

yes!!!!

OpenStudy (anonymous):

mk, the answer is pretty simple(if im doing this correctly) give me a min to write it out in equation form

OpenStudy (anonymous):

\[1/2 \log_{8} x = 3\log_{8} 3\] whic\[x = \sqrt{9}\]h becomes \[\log_{8} \sqrt{x} = \log_{8} 3^{3}\] since any value to the exponent 1/2 is a square root. now drop the log 8 and it gives you \[\sqrt{x} = 9\] where \[x = \sqrt{9}\] therefore, x = 3

OpenStudy (anonymous):

wow my whole message got fluttered up. hold on

OpenStudy (anonymous):

k well im not gonna rewrite it ignore the square root 9 at the top. what u need ot grasp is by following logarithm rules, you know that the values in front of the logaritms become exponent of values after log. therefore x ^ 1/2 is simply square root of x and on the other side 3^3 is 9. actually i think i made a mistake, when u get rid of the sqrt root u square both sides, so x is actually equal to 9^2, which is 81.

OpenStudy (anonymous):

hope it helps.

OpenStudy (anonymous):

the answer was 729 ?! oh well i alreaody its already graded now! thanks for your help tho!

OpenStudy (anonymous):

oh pellet... flutter dude 3^3 is 27 MY BAD! so 27 ^2 is 729!!!

OpenStudy (anonymous):

so it was the right way to do it i just did 3^3 wrong :(...im sorry!